Hassett, B. et al. () “Balanced line bundles and equivariant compactifications of homogeneous spaces,” International Mathematics Research Notices, Vol. , Article ID , 33 pages. doi:10.1093/imrn/

Balanced line bundles and equivariant compactifications of homogeneous spaces

Brendan Hassett^{1}, Sho Tanimoto^{1}, and Yuri Tschinkel^{2}^{3}

1Department of Mathematics, Rice University MS 136, Houston, Texas 77251-1892 USA, ^{2}Courant Institute, New York University, New York, NY 10012 USA, and ^{3}Simons Foundation, 160 Fifth Av., New York, NY 10010

Correspondence to be sent to: hassett@rice.edu, sho.tanimoto@rice.edu, tschinkel@cims.nyu.edu

Manin’s conjecture predicts an asymptotic formula for the number of rational points of bounded height on a smooth projective variety X in terms of global geometric invariants of X. The strongest form of the conjecture implies certain inequalities among geometric invariants of X and of its subvarieties. We provide a general geometric framework explaining these phenomena, via the notion of balanced line bundles, and prove the required inequalities for a large class of equivariant compactifications of homogeneous spaces.

1 Introduction

Let X be a smooth projective variety over a number field. It is generally hoped that global geometric properties of X should be reflected in its arithmetic properties. For instance, assume that its anticanonical class −K_{X} is ample. It has been conjectured that such X satisfy:

Potential Density: there exists a finite extension F of the ground field such that X(F) is Zariski dense (see [HT00a], [BT99] for first results in this direction and [Cam04], [Abr09] for a description of a general framework).

Supposing that X has dense rational points over F, we can ask for quantitative versions of density:

Asymptotic Formulas: Let L = (L, ∥ · ∥) be an ample, adelically metrized, line bundle on X and H_{L} the associated height (for definitions and more background see, e.g., [Tsc09, Section 4.8]). Then there exists a Zariski open X^{∘} ⊂ X such that

#{x ∈ X^{∘}(F) | H_{L}(x) ≤ B} ∼ c(X, L)B^{a(X,L)} log(B)^{b(X,L)−1},

Received ; Revised ; Accepted Communicated by A. Editor

\begin{matrix}c \\ ◯ \end{matrix} The Author . Published by Oxford University Press. All rights reserved. For permissions, please e-mail: journals.permissions@oxfordjournals.org.

2 B. Hassett et al.

as B → ∞. Here a(X, L) and b(X, L) are certain geometric constants introduced in this context in

[FMT89] and [BM90] (and recalled in Section 2) and c(X, L) is a Tamagawa-type number defined in [Pey95], [BT98b].

When L = −K_{X} the main term of the asymptotic formula reads

#{x ∈ X^{∘}(F) | H_{−K_{X}} (x) ≤ B} ∼ c(X, −K_{X})B log(B)^{rk Pic(X)−1}

as B → ∞, where −K_{X} is the metrized anticanonical bundle. For a survey addressing both aspects and containing extensive references, see [Tsc09].

The Asymptotic Formulas raise many formal questions. How do we choose X^{∘} ⊂ X? Clearly, we want to exclude subvarieties Y ⊊ X contributing excessively to the number of rational points. For example, if X is a split cubic surface and L = −K_{X} then lines on X contribute on the order of B^{2} points of height ≤ B, more than the B log(B)^{6} points expected from X^{∘}.

Furthermore, we should consider carefully whether to include subvarieties Y ⊊ X contributing rational points at the same rate as those from X^{∘}. For example, if X ⊂ ℙ^{5} is a complete intersection of two quadrics then each line of X contributes on the order of B^{2} points, the same as the conjectured total for X^{∘} (see Example 3.11). These lines are parametrized by an abelian surface. Including such subvarieties must have implications for the interpretation of the Tamagawa-type constant.

Returning to the case of general L, in order for the Asymptotic Formula to be internally consistent, all Y ⊊ X meeting X^{∘} must satisfy

(a(Y, L|_{Y}), b(Y, L|_{Y})) ≤ (a(X, L), b(X, L))

in the lexicographic order. Moreover, if the constant c(X, L) is to be independent of the open set X^{∘} ⊂ X we must have

(a(Y, L|_{Y})), b(Y, L|_{Y})) < (a(X, L), b(X, L)).  (1)

However, there exist varieties of dimension ≥ 3 where these properties fail; these provide counterexamples to the Asymptotic Formulas [BT96b]. On the other hand, no counterexamples are known in the equivariant context, when X is an equivariant compactification of a linear algebraic group G or of a homogeneous space H\G, and asymptotic formulas for the number of points of bounded height have been established for many classes of such compactifications (see [Tsc09]).

These arithmetic considerations motivate us to introduce and study the notion of balanced line bundles. A balanced line bundle is one for which general subvarieties Y satisfy the inequalities above, yielding compatibility with Manin’s predictions. In this paper, we establish basic properties of balanced line bundles and investigate varieties that carry such line bundles.

Balanced line bundles 3

Throughout the paper, we assume that the ground field is an algebraically closed field of characteristic zero. In Section 2, we introduce geometric invariants a(X, L) and b(X, L) and explore their basic propeties, i.e., questions of rationality and birational invariance. Then in Section 3, we define the notion of balanced line bundles and study some important varieties including flag varieties and some smooth Fano 3-folds. One important observation is

Proposition 1.1. Let X be a generalized flag variety and L a big line bundle on X. Then L is balanced with respect to any smooth variety if and only if L is proportional to −K_{X}.

In Section 4, we study balanced line bundles on del Pezzo surfaces. We obtain the complete characterization of balanced line bundles which is described below:

Proposition 1.2. Let X be a del Pezzo surface and L a big Cartier divisor on X. Then L is balanced if and only if the ℚ-divisor a(X, L)L + K_{X} is linearly equivalent to a rigid effective divisor on X.

Here we obtain an interesting equivalence between an analytic concept and a geometric concept. It claims that the rigidity of the adjoint divisor and the balanced property are equivalent. This observation has been made already in Proposition 1.1. In Section 5, we study the geometry of equivariant compactifications of homogeneous spaces and obtain the following result, which is a main result of this paper:

Theorem 1.3. Let

H ⊂ M ⊂ G

be connected linear algebraic groups. Let X be a smooth projective G-equivariant compactification of H\G and Y ⊂ X the induced compactification of H\M. Assume that the projection G → M\G admits a rational section. Then −K_{X} is balanced with respect to Y, i.e., inequality (1) holds for L = −K_{X}.

A version of this geometric result, for G = 𝔾_{a}^{n}, appeared in [CLT02, Section 7], where it was used to bound contributions from nontrivial characters to the Fourier expansion of height zeta functions and, ultimately, to prove asymptotic formulas for the number of rational points of bounded height (Manin’s conjecture) for equivariant compactifications of 𝔾_{a}^{n}. Another application can be found in [GTBT11], where this theorem plays an important role in an implementation of ideas from ergodic theory (mixing) in a proof of Manin’s conjecture for equivariant compactifications of G\G^{n}, where G is an absolutely simple linear algebraic group, acting diagonally on G^{n}. In Section 6 we investigate balanced line bundles on toric varieties in the context of the Minimal Model Program. The main result is

Proposition 1.4. Let X be a smooth projective toric variety and L a big Cartier divisor. Then L is balanced with respect to all subtoric varieties if and only if the ℚ-divisor a(X, L)L + K_{X} is linearly equivalent to a rigid effective divisor on X.

Again, we observe the equivalence between the balanced property and the rigidity of the adjoint divisor.

4 B. Hassett et al.

2 Generalities

Definition 2.1. Let V be a finite dimensional vector space over ℝ. A closed convex cone Λ ⊂ V is a closed subset which is closed under linear combinations with non-negative real coefficients. An extremal face F ⊂ Λ is a closed convex subcone of Λ such that if u, v ∈ Λ and u + v ∈ F then u, v ∈ F. A supporting function is a linear functional σ : V → ℝ such that σ ≥ 0 on Λ. A face of the form

F′ = {σ = 0} ∩ Λ

is called a supported face. A supported face is an extremal face, but the converse is not true, in general. The converse does hold when Λ is locally finitely generated in a neighborhood of F, i.e., there exist finitely many linear functionals

λ_{i} : V → ℝ

such that λ_{i}(v) > 0 for v ∈ F \ {0} and

Λ ∩ {v : λ_{i}(v) ≥ 0 for any i},

is finitely generated. Note that when Λ is strict, i.e., does not contain a line, then {0} is a supported face.

We work over an algebraically closed field of characteristic zero. A variety is an integral separated scheme over this field. Let X be a smooth projective variety. We use

Λ_{eff}(X) ⊂ NS(X, ℝ) ⊂ H^{2}(X, ℝ)

to denote the pseudo-effective cone, i.e., the closure of effective ℚ-divisors on X in the real Néron-Severi group NS(X, ℝ). Another common notation in the literature is \overline{NE}^{1}(X). Note that the pseudo-effective cone is strict and convex [BFJ09, Prop. 1.3]. Let Λ_{eff}^{∘}(X) denote the interior of the pseudo-effective cone; a divisor class D on X is big if [D] ∈ Λ_{eff}^{∘}(X). We denote the dual cone of the cone of pseudo-effective divisors by \overline{NM}_{1}(X). This is the closure of the cone generated by movable curves ([BDPP13].)

A rigid effective divisor is a reduced divisor D ⊂ X such that

H^{0}(O_{X}(nD)) = 1 ∀n ≥ 1.

If D is rigid with irreducible components D_{1}, …, D_{r} then

H^{0}(O_{X}(n_{1}D_{1} + … + n_{r}D_{r})) = 1 ∀n_{1}, …, n_{r} ≥ 1

Balanced line bundles 5

and

span(D_{1}, …, D_{r}) ∩ Λ_{eff}^{∘}(X) = ∅.  (2)

Definition 2.2. Assume that L is a big Cartier divisor on X. The Fujita invariant is defined by

a(X, L) = inf{a ∈ ℝ : aL + K_{X} ∼ an effective ℝ-divisor }
= min{a ∈ ℝ : a[L] + [K_{X}] ∈ Λ_{eff}(X)}.

Remark 2.3. Note that a(X, L) only depends on the numerical class of L.

Note that the Fujita invariant is positive if and only if K_{X} is not pseudo-effective. The invariant

κϵ(X, L) = −a(X, L)

was introduced and studied by Fujita under the name Kodaira energy [Fuj97] (see also [Fuj87], [Fuj92], [Fuj96]). A similar invariant

σ(X, L) = dim(X) + 1 − a(X, L)

appeared in [Som86] under the name spectral value.

Remark 2.4. A smooth projective variety X is uniruled if and only if K_{X} is not pseudo-effective [BDPP13], [Laz04b, Cor. 11.4.20].

The following result was conjectured by Fujita and proved by Batyrev for threefolds and [BCHM10, Cor. 1.1.7] in general. (See [DC12, Corollay 3.7 and Corollary 3.8] for recent generalizations.)

Theorem 2.5. Let X be projective with Kawamata log terminal singularities such that K_{X} is not pseudoeffective, and L an ample Cartier divisor on X. Then a(X, L) is rational.

However, this property can fail when L is big but not ample, and the following example going back to Cutkosky [Cut86] was suggested to us by Brian Lehmann:

Example 2.6. [Leh12, Example 4.9] Let Y be an abelian surface with Picard rank at least 3. The cone of nef divisors and the cone of pseudo-effective divisors coincide, and the boundary of these cones is circular. Let N be a Cartier divisor on Y such that −N is ample and X := ℙ(O ⊕ O(N)). Let π : X → Y denote the projection morphism and S ⊂ X the section corresponding to the quotient map O ⊕ O(N) → O(N). Every divisor on X is linearly equivalent to tS + π^{∗}D where D is a divisor on Y. In particular, K_{X} is linearly equivalent to −2S + π^{∗}N. The cone of pseudo-effective divisors Λ_{eff}(X) is generated by S and π^{∗}Λ_{eff}(Y).

6 B. Hassett et al.

Consider a big ℚ-divisor L = tS + π^{∗}D, where t > 0 and D is a big ℚ-divisor on Y. If t is sufficiently large, then a(X, L) is characterized by the condition a[D] + [N] ∈ ∂Λ_{eff}(Y). However, the boundary of Λ_{eff}(Y) is circular, and a(X, L) ∉ ℚ, in general.

From the point of view of Manin’s conjecture, the global geometric invariants involved in its formulation should be functorial for birational transformations, and indeed this holds for the Fujita invariant: Proposition 2.7. Let β : X̃ → X be a birational morphism of projective varieties, where X̃ is smooth, and X is normal and has only canonical singularities. Assume K_{X} is not pseudo-effective and L is a Cartier divisor which is big. Setting L̃ = β^{∗}L, we have

a(X, L) = a(X̃, L̃).

Proof. Since X has canonical singularities, we have

KX̃ = β^{∗}K_{X} + ∑_{i} d_{i}E_{i},

where the E_{i} are the irreducible exceptional divisors and the d_{i} are nonnegative rational numbers. It follows that for sufficiently divisible integers m, n ≥ 0 we have

Γ(O_{X}(mK_{X} + nL)) = Γ(OX̃(mKX̃ − ∑_{i} md_{i}E_{i} + nL̃))
= Γ(OX̃(mKX̃ + nL̃)),

where the second equality reflects the fact that allowing poles in the exceptional locus does not increase the number of global sections. In particular, effective divisors supported in the exceptional locus of β are rigid. It follows from the assumption that no multiple of KX̃ is effective and that a(X̃, L̃) ≥ 0. Definition 2.2 gives a(X̃, L̃) = a(X, L) > 0.

Next, we discuss the second geometric invariant appearing in Manin’s conjecture. Definition 2.8. Let X be a projective variety with only ℚ-factorial terminal singularities such that K_{X} is not pseudo-effective. Let L be a big Cartier divisor on X. Define

b(X, L) = the codimension of the minimal supported face of
Λ_{eff}(X) containing the ℝ-divisor a(X, L)L + K_{X}.

Remark 2.9. Again b(X, L) only depends on the numerical class of L

Balanced line bundles 7

This definition is relatively easy to grasp when Λ_{eff}(X) is finitely generated, which holds in a number of cases:

• A projective variety X is log Fano if there exists an effective ℚ-divisor Δ on X such that (X, Δ) is divisorially log terminal (see [BCHM10, p. 424]) and −(K_{X} + Δ) is ample. If X is log Fano then the Cox ring of X is finitely generated [BCHM10, Cor. 1.3.2], so in particular, Λ_{eff}(X) is finite rational polyhedral and generated by effective divisors.

• Let X be a smooth projective variety that is toric or an equivariant compactification of the additive group 𝔾_{a}^{n}. Then Λ_{eff}(X) is generated by boundary divisors, i.e., irreducible components of the complement of the open orbit, [HT99, Thm. 2.5], [BT95, Prop.1.2.11].

Let X be a smooth projective variety with Λ_{eff}(X) generated by a finite number of effective divisors and Pic(X)_{ℚ} = NS(X, ℚ). Since each irreducible rigid effective divisor on X is a generator of Λ_{eff}(X) (cf. (2)), we have

Z := ∪_{rigid effective}D

is a Zariski closed proper subset of X.

One of the reasons for adopting the terminology of supported faces in the definition of b(X, L) is to simplify the verification of its birational invariance:

Proposition 2.10. Let X be a normal ℚ-factorial terminal projective variety such that K_{X} is not pseudoeffective and β : X̃ → X a smooth resolution. Let L be a big Cartier divisor on X and put L̃ = β^{∗}L. Then

b(X, L) = b(X̃, L̃).

Proof. Let F be the minimal supported face of Λ_{eff}(X) containing an ℝ-divisor a(X, L)L + K_{X} and F̃ be the minimal supported face of Λ_{eff}(X̃) containing a(X̃, L̃)L̃ + KX̃. The vector spaces generated by F and F̃ will be denoted by V_{F} and VF̃, respectively. There exists a nef cycle ξ ∈ \overline{NM}_{1}(X̃) such that

F̃ = {ξ = 0} ∩ Λ_{eff}(X̃).

Let E_{1}, …, E_{n} be irreducible components of the exceptional locus of β. The Negativity Lemma ([BCHM10, Lemma 3.6.2]) implies that NS(X̃) is a direct sum of β^{∗}NS(X) and ℝ[E_{i}]’s. Since X has only terminal singularities, it follows from Proposition 2.7 that

a(X̃, L̃)L̃ + KX̃ = β^{∗}(a(X, L)L + K_{X}) + ∑_{i} d_{i}E_{i},

8 B. Hassett et al.

where the d_{i}’s are positive rational numbers. This implies that F̃ contains β^{∗}(a(X, L)L + K_{X}) and the E_{i}’s. Thus a(X, L)L + K_{X} is contained in an extremal face supported by a supporting function β_{∗}ξ, i.e.,

{β_{∗}ξ = 0} ∩ Λ_{eff}(X),

so this supported face also contains F. We get a well-defined injection

Φ : V_{F} ↪ VF̃/(∑_{i} ℝE_{i}).

On the other hand, let η ∈ \overline{NM}_{1}(X) be a nef cycle supporting F. Consider a linear functional η̃: NS(X̃) → ℝ defined by

η̃ ≡ η on β^{∗}NS(X) and η̃ · E_{i} = 0 for any i.

The projection from NS(X̃) to β^{∗}NS(X) maps pseudo-effective divisors to pseudo-effective divisors so that η̃ ∈ \overline{NM}_{1}(X̃). Moreover,

η̃ · (a(X̃, L̃)L̃ + KX̃) = η̃ · (β^{∗}(a(X, L)L + K_{X}) + ∑_{i} d_{i}E_{i}) = 0,

so that {η̃ = 0} ∩ Λ_{eff}(X̃) contains F̃. It follows that Φ is bijective and our assertion is proved.

Definition 2.11. Let X be a uniruled projective variety with a big Cartier divisor L. We define

a(X, L) = a(X̃, β^{∗}L), b(X, L) = b(X̃, β^{∗}L)

where β : X̃ → X is some resolution of singularities.

Note that KX̃ is not pseudo-effective by Remark 2.4; Propositions 2.7 and 2.10 guarantee the invariants are independent of the choice of resolution.

For the anticanonical line bundle, invariants are computed as follows:

Example 2.12 (The anticanonical line bundle). Let X be a projective variety with only ℚ-factorial terminal singularities. As in the smooth case, the cone Λ_{eff}(X) of pseudo-effective divisors is strict. When the anticanonical class −K_{X} is big, we have

a(X, −K_{X}) = 1, b(X, −K_{X}) = rk NS(X).

Let β : X̃ → X be a smooth resolution and E_{1}, …, E_{n} the irreducible components of the exceptional locus; we have

KX̃ = β^{∗}K_{X} + ∑_{i} d_{i}E_{i},

Balanced line bundles 9

where d_{i} ∈ ℚ_{>0}, for all i. Hence the minimal extremal face containing −β^{∗}K_{X} + KX̃ contains a simplicial cone F := ⊕_{i}ℝ_{≥0}[E_{i}]. The fact that F is a supported face follows from [Bou04, Theorem 3.19], which asserts that the pseudo-effective cone is locally polyhedral in this region, generated by the prime exceptional divisors. Hence we may compute

b(X̃, −β^{∗}K_{X}) = rk NS(X̃, ℝ) − n = rk NS(X, ℝ) = b(X, −K_{X}).

Remark 2.13. There exist projective bundles over curves of arbitrary genus g > 0 with big anticanonical divisor [KMM92, 3.13]. When g > 1 these cannot have potentially dense rational points.

Example 2.14. Example 2.12 can be generalized as follows: Let X be a smooth projective variety and D = ∑_{i} e_{i}E_{i} an effective ℝ-divisor whose numerical dimension ν(D) is zero (see [Nak04, Section V.2] or [Leh13, Theorem 1.1] for definitions). The minimal extremal face containing D contains F = ⊕_{i}ℝ_{≥0}[E_{i}], and we claim that F is a supported face of Λ_{eff}(X). First we prove that F is an extremal face. Let u, v ∈ Λ_{eff}(X) such that u + v ∈ F. For any pseudo-effective numerical class α, we denote the negative part and the positive part of the divisorial Zariski decomposition of α by N_{σ}(α) and P_{σ}(α) = α − [N_{σ}(α)] ∈ Λ_{eff}(X) respectively (see [Nak04, Section III.1], also [Leh13, Section 3]). The assumption ν(D) = 0 implies that

u + v ≡ N_{σ}(u + v) ≤ N_{σ}(u) + N_{σ}(v).

This implies that N_{σ}(u) ≡ u and N_{σ}(v) ≡ v so that u, v ∈ F. Again by [Bou04, Theorem 3.19], the cone Λ_{eff}(X) is locally rational polyhedral in a neighborhood of F. Hence F is a supported face.

Little is known about the geometric meaning of the invariant b(X, L), in general. Here we consider situations relevant for our applications to equivariant compactifications of homogeneous spaces.

Definition 2.15. Let X be a ℚ-factorial terminal and projective variety and D an ℝ-divisor in the boundary of Λ_{eff}(X). We say D is locally rational polyhedral if either D ≡ 0 numerically or there exist finitely many linear functionals

λ_{i} : NS(X, ℚ) → ℚ

such that λ_{i}(D) > 0 and

Λ_{eff}(X) ∩ {v : λ_{i}(v) ≥ 0 for any i},

is finite rational polyhedral and generated by effective ℚ-divisors. In this case, the minimal extremal face F containing D is supported by a supporting function.

Theorem 2.16. Let X be a ℚ-factorial terminal projective variety such that K_{X} is not pseudo-effective and L a big Cartier divisor on X. Suppose that a(X, L)L + K_{X} has the form c(A + K_{X} + Δ), where A is an ample

10 B. Hassett et al.

ℝ-divisor, (X, Δ) a Kawamata log terminal pair, and c > 0. Then a(X, L)L + K_{X} is locally rational polyhedral and a(X, L) is rational.

Proof. When D = a(X, L)L + K_{X} ≡ 0, then a(X, L) is rational. Suppose that D ≢ 0. The local finiteness of the pseudo-effective boundary is proved in [Leh12, Proposition 3.3] by using the finiteness of the ample models [BCHM10, Corollary 1.1.5]. Moreover, Lehmann proved that the pseudo-effective boundary is locally defined by movable curves. The generation by effective ℚ-divisors follows from the non-vanishing theorem [BCHM10, Theorem D].

In particular, if L is ample then a(X, L)L + K_{X} is locally rational polyhedral. We have already seen in Example 2.6 that the local finiteness is no longer true if we only assume that L is big. However, there are certain cases where the local finiteness of a(X, L)L + K_{X} still holds for any big Cartier divisor L:

Example 2.17 (Surfaces). Let X be a smooth projective surface such that K_{X} is not pseudo-effective. Let L be a big Cartier divisor on X. If D = a(X, L)L + K_{X} is numerically trivial, then a(X, L) is rational. Suppose that D is a non-zero pseudo-effective divisor. We prove that D is locally rational polyhedral. We consider the Zariski decomposition of D = P + N, where P is a nef ℝ-divisor and N is the negative part of D. The boundary of Λ_{eff}(X) is locally rational polyhedral away from the nef cone (see [Bou04, Theorem 3.19] and [Bou04, Theorem 4.1]). Thus, if N is non-zero, then our assertion follows. Suppose that N is zero. Since a(X, L)L · P + K_{X} · P = D · P = 0 and L · P > 0, we have K_{X} · P < 0. Therefore, D = P is sitting in the K_{X}- negative part of the boundary of the cone of curves. Now our assertion follows from Mori’s cone theorem. In particular, a(X, L) is a rational number.

In general, a(X, L)L + K_{X} is not necessarily of the form of the adjoint divisor c(A + K_{X} + Δ), where A is an ample ℝ-divisor, (X, Δ) a Kawamata log terminal pair, and c > 0.

Example 2.18 (Equivariant compactifications of the additive groups). Let X be a smooth projective equivariant compactification of the additive group 𝔾_{a}^{n}. Then Λ_{eff}(X) is a simplicial cone generated by boundary components, by [HT99, Theorem 2.5]. However, this cannot be explained from Theorem 2.16. Indeed, consider the standard embedding of 𝔾_{a}^{3} into ℙ^{3}:

𝔾_{a}^{3} ∋ (x, y, z) ↦ (x : y : z : 1) ∈ ℙ^{3}.

This is an equivariant compactification, and the group action fixes every point on the boundary divisor D, a hyperplane section. Let X be an equivariant blow up of 12 generic points on a smooth cubic curve in D. Write H for the pullback of the hyperplane class and E_{1}, …, E_{12} for the exceptional divisors. Consider

L = 4H − E_{1} − ⋯ − E_{12}.

Balanced line bundles 11

Then L is big and nef, but not semi-ample (see [Laz04a, Section 2.3.A] for more details). In particular, the section ring of L is not finitely generated. On the other hand, consider

Λ_{adj}(X) = {Γ ∈ Λ_{eff}(X) | Γ = c(A + K_{X} + Δ)},

where A is an ample ℝ-divisor, (X, Δ) a Kawamata log terminal pair, and c a positive number. Then Λ_{adj}(X) forms a convex cone. The existence of non-finitely generated divisors and [BCHM10, Corollary 1.1.9] imply that Λ_{adj}(X) ⫋ Λ_{eff}(X).

It is natural to expect that the invariant b(X, L) is related to the canonical fibration associated to a(X, L)L + K_{X}. A sample result in this direction is:

Proposition 2.19. Let X be a smooth projective variety such that K_{X} is not pseudo-effective. Let L be a big Cartier divisor and assume that D = a(X, L)L + K_{X} is locally rational polyhedral and semi-ample. Let π : X → Y be the semi-ample fibration of D. Then

b(X, L) = rk NS(X) − rk NS_{π}(X),

where NS_{π}(X) is the lattice generated by π-vertical divisors, i.e., divisors M ⊂ X such that π(M) ⊊ Y

Proof. Let F be the minimal extremal face of Λ_{eff}(X) containing D = a(X, L)L + K_{X} and V_{F} the vector space generated by F. We claim that V_{F} = NS_{π}(X). Let H be an ample ℚ-divisor on Y such that π^{∗}H = D. Let M be a π-vertical divisor on X. Then for sufficiently large m, there exists an effective Cartier divisor H′ such that mH ∼ H′ and the support of H′ contains π(M). Thus mD = mπ^{∗}H ∼ π^{∗}H′ ∈ F and the support of π^{∗}H′ contains M. We conclude that M ∈ F, and this proves that NS_{π}(X) ⊂ V_{F}. Next, let X_{y} be a general fiber of π and C ⊂ X_{y} a movable curve on X such that [C] is in the interior of \overline{NM}_{1}(X_{y}). Then

F_{C} = {[C] = 0} ∩ Λ_{eff}(X),

is an extremal face containing D. The minimality implies F ⊂ F_{C}. On the other hand, the local rational finiteness of D implies that there exist effective ℚ-divisors D_{1}, …, D_{n} ∈ F such that D_{1}, …, D_{n} form a basis of V_{F}. Since D_{1} · C = ⋯ = D_{n} · C = 0, the supports of D_{i}’s are π-vertical. Hence it follows that V_{F} ⊂ NS_{π}(X).

Remark 2.20. When L is ample, it follows from [KMM87, Lemma 3.2.5] that the codimension of the minimal extremal face of nef cone containing D is equal to the relative Picard rank ρ(X/Y).

In Section 6, we explore this further in the case of toric varieties.

12 B. Hassett et al.

3 Balanced line bundles

Definition 3.1. Let X be a uniruled projective variety, L a big line bundle on X, and Y ⊊ X an irreducible uniruled subvariety. L is weakly balanced with respect to Y if

• L|_{Y} is big;

• a(Y, L|_{Y}) ≤ a(X, L);

• if a(Y, L|_{Y}) = a(X, L) then b(Y, L|_{Y}) ≤ b(X, L).

It is balanced with respect to Y if it is weakly balanced and one of the two inequalities is strict.

L is weakly balanced (resp. balanced) on X if there exists a Zariski closed subset Z ⊊ X such that L is weakly balanced (resp. balanced) with respect to every Y not contained in Z. The subset Z will be called exceptional.

Remark 3.2. The restriction to uniruled subvarieties is quite natural: If Y is a smooth projective variety that is not uniruled and Y → X is a morphism such that L|_{Y} is big, then a(Y, L|_{Y}) ≤ 0 < a(X, L) (see Remark 2.4).

Remark 3.3. Any big Cartier divisor can be expressed as a sum of an ample ℚ-divisor A and an effective ℚ-divisor E. In particular, L|_{Y} is big for any Y ⊈ Supp(E).

We first explore these properties for projective homogeneous spaces:

Proposition 3.4. Let G be a connected semi-simple algebraic group, P ⊂ G a parabolic subgroup and X = P \G the associated generalized flag variety. Let L be a big Cartier divisor on X. We have:

• if L is not proportional to −K_{X} then L is not balanced but is weakly balanced with respect to smooth subvarieties Y ⊂ X;

• if L is proportional to −K_{X} then L is balanced with respect to smooth subvarieties.

Proof. For generalized flag varieties, the nef cone and the pseudo-effective cone coincide so that L is ample. Moreover, the nef cone of a flag variety is finitely generated by semi-ample line bundles. Also note that since every rationally connected smooth proper variety is simply connected, all parabolic subgroups are connected.

Assume that L is not proportional to the anticanonical bundle, i.e., D = a(X, L)L + K_{X} is a non-zero effective ℚ-divisor. Let π : X → X′ be the semi-ample fibration of D. Then X′ is also a G-variety so that there exists a parabolic subgroup P′ ⊃ P such that X′ = P′\G and π is the natural projection map. We have the following exact sequence:

0 → Pic(P′\G)_{ℚ} → Pic(P \G)_{ℚ} → Pic(P \P′)_{ℚ} → 0.

Indeed, the surjectivity follows from [KKV89, Proposition 3.2(i)]. Note that the picard group of a linear algebraic group is necessary finite (see [KKLV89, Proposition 4.5]). Then the exactness of other parts follows from

Balanced line bundles 13

[KMM87, Lemma 3.2.5]. Let W be a fiber of π. Then a(X, L) = a(W, L|_{W}) since K_{X}|_{W} = K_{W} and D|_{W} = 0. The exact sequence and Remark 2.20 imply that

b(X, L) = ρ(X/X′) = rk Pic(W) = b(W, L|_{W}).

Thus L is not balanced with respect to any fiber of π.

Let L be an arbitrary ample Cartier divisor and Y ⊂ X a smooth subvariety. Let 𝔤 be the Lie algebra of G. For any ∂ ∈ 𝔤, we can construct a global vector field ∂^{X} on X such that for any open set U ⊂ X and any f ∈ O_{X}(U),

∂^{X}(f)(x) = ∂_{g}f(x · g)|_{g=1}.

It follows that the normal bundle N_{Y/X} is globally generated and its determinant is as well.

The restriction

aL|_{Y} + K_{Y} = (aL + K_{X})|_{Y} + det(N_{Y/X})

with det(N_{Y/X}) globally generated hence contained in Λ_{eff}(Y). Thus we have a(Y, L|_{Y}) ≤ a(X, L). Suppose equality holds; our goal then is to prove that

b(Y, L|_{Y}) ≤ b(X, L),

and the strict inequality holds when L is proportional to −K_{X}. Let D = aL + K_{X} which is semi-ample.

First we assume that det(N_{Y/X}) and D|_{Y} are trivial so that N_{Y/X} is the trivial vector bundle of rank r = codim(Y, X). The above construction of vector fields defines a surjective map:

φ : 𝔤 → H^{0}(Y, N_{Y/X}).

We may assume that e = P ∈ Y so that the Lie algebra 𝔭 of P is contained in the kernel of φ. Consider the Hilbert scheme Hilb(X) and note that H^{0}(Y, N_{Y/X}) is naturally isomorphic to the Zariski tangent space of Hilb(X) at [Y]. Consider the morphism:

π : G ∋ g ↦ [Y · g] ∈ Hilb(X).

Since Y is Fano, H^{1}(Y, N_{Y/X}) = 0, Hilb(X) is smooth at [Y], and

dim_{[Y]}(Hilb(X)) = r.

Moreover, since φ is surjective, π is a smooth morphism and π(G) is a smooth open subscheme in Hilb(X). Let

14 B. Hassett et al.

H be the connected component of Hilb(X) containing [Y] and P′ = Stab(Y). Since the kernel of φ contains 𝔭, we have P ⊂ P′. This implies that π(G) = P′\G is open and closed so that

H = π(G) = P′\G.

In particular, dim(G) − dim(P′) = r, so the kernel of φ is exactly equal to the Lie algebra 𝔭′ of P′. Consider the universal family U ⊂ X × H on H. It follows that G acts on U transitively, and we conclude that U = P \G and Y = P \P′. Since D|_{Y} is trivial, b(Y, L|_{Y}) = rk Pic(P \P′). The exact sequence, which we discussed before, and [KMM87, Lemma 3.2.5] indicate that rk Pic(P \P′) = ρ(X/H), the relative Picard rank of π. It follows from Remark 2.20 and the triviality of D|_{Y} that

ρ(X/H) ≤ b(X, L).

When L is proportional to −K_{X}, b(X, L) = rk NS(X) so that the strict inequality holds.

In the general case, we still know that N_{Y/X} and its determinant are globally generated. Consider the semi-ample fibration Y → W associated to aL|_{Y} + K_{Y} = D|_{Y} + det(N_{Y/X}) with generic fiber Y_{w}, which is smooth. Note that N_{Y_{w}/X} is trivial, as det(N_{Y/X})|_{Y_{w}} and N_{Y_{w}/Y} are both trivial. D|_{Y} is also trivial. The above construction shows that Y_{w} is the fiber of a fibration ρ : X → B with W ⊂ B; Y is the pullback of W. Theorem 2.16 and Proposition 2.19 imply that

b(Y, L|_{Y}) = rk NS(Y) − rk NS_{ρ}(Y).

The restriction map

Φ : NS(Y)/NS_{ρ}(Y) → NS(Y_{w}),

is injective; this follows from [KMM87, Lemma 3.2.5]. Note that ρ is an isotrivial family. Hence we have

b(Y, L|_{Y}) ≤ b(Y_{w}, L|_{Y_{w}}) ≤ b(X, L).

If L is proportional to −K_{X}, then the last inequality is a strict inequality.

We can also analyze the balanced condition with respect to hypersurfaces: Proposition 3.5. Let X be a smooth Fano variety of Picard rank one and Y ⊂ X an irreducible smooth effective divisor. Then −K_{X} is balanced with respect to Y.

Proof. For smooth divisors Y ⊂ X the claim follows from adjunction formula:

−K_{X}|_{Y} + K_{Y} ∈ Λ_{eff}(Y)^{∘},

Balanced line bundles 15

because Y is an ample divisor on X. Thus we obtain

a(Y, −K_{X}|_{Y}) < a(X, −K_{X}) = 1.

However, this may fail when Y is singular:

Example 3.6 (Mukai-Umemura 3-folds, [MU83]). Consider the standard action of SL_{2} on V = ℂx ⊕ ℂy. Let R_{12} = Sym^{12}(V) be a space of homogeneous polynomials of degree 12 in two variables and f ∈ R a form with distinct roots. Let X be the Zariski closure of the SL_{2}-orbit SL_{2} · [f] ⊂ ℙ(R_{12}). Then X is a smooth Fano 3-fold of index 1 with Pic(X) = ℤ, for some special f. The complement of the open orbit SL_{2} · [f] is an irreducible divisor

D = \overline{SL_{2} · [x^{11}y]} = SL_{2} · [x^{11}y] ∪ SL_{2} · [x^{12}],

a hyperplane section on ℙ(R_{12}) whose class generates Pic(X). Furthermore, D is the image of ℙ^{1} × ℙ^{1} by a linear series of bidegree (11, 1), which is injective, an open immersion outside of the diagonal, but not along the diagonal. In particular, D is singular along the diagonal.

Let β : D̃ → D be the normalization of D which is isomorphic to ℙ^{1} × ℙ^{1}. Then −β^{∗}K_{X}|D̃ is a line bundle of bidegree (11, 1) so that

a(D, −K_{X}|_{D}) = a(D̃, −β^{∗}K_{X}|D̃) = 2 > 1 = a(X, −K_{X}).

Thus Proposition 3.5 does not hold for D.

Remark 3.7. The authors do not know whether Proposition 3.5 holds for singular surfaces in ℙ^{3}. It is quite interesting to see whether the balanced property holds for very singular rational surfaces in ℙ^{3}.

For Fano varieties of index one, one might hope to use the Fujita invariant a(X, L) to identify the exceptional locus X \ X^{∘}. However, this is quite non-trivial even in the following situation, considered in [Deb03] (see also [LT10] and [Beh06]):

Conjecture 3.8 (Debarre - de Jong conjecture). Let X ⊂ ℙ^{n} be a Fano hypersurface of degree d ≤ n. Then the dimension of the variety of lines is 2n − d − 3. In particular, when d = n, for any line C, we have

a(C, −K_{X}|_{C}) = 2 > 1 = a(X, −K_{X}).

The conjecture predicts that the dimension of the variety of lines is n − 3 so that lines will not sweep out X.

Some of simplest examples of Fano threefolds fail to be balanced.

16 B. Hassett et al.

Example 3.9. Let X ⊂ ℙ^{4} be a smooth cubic threefold, which is Fano of index 2. The Picard group of X is generated by the hyperplane class L. By Proposition 3.5, −K_{X} is balanced with respect to every smooth divisor on X. Let Y ⊂ X be a line. Note that 2L restricts to the anticanonical class on X and Y, and b(Y, L|_{Y}) = b(X, L) = 1. Thus −K_{X} is weakly balanced, but not balanced, with respect to Y. Since the family of lines dominates X, −K_{X} is not balanced on X. Remark 3.10. However, assume that X is defined over a number field. The family of lines dominating X are surfaces of general type, which embeded into their Albanese varieties. By Faltings’ theorem, lines defined over a fixed number field lie on a proper subvariety and cannot dominate X.

One can rephrase the above example in the following way: the anticanonical line bundle of a smooth cubic threefold is not geometrically balanced, but arithmetically balanced. Next example shows that it is also possible that the anticanonical line bundle is not even arithmetically balanced. Example 3.11. Let X ⊂ ℙ^{5} denote a smooth complete intersection of two quadrics. The anticanonical class −K_{X} = 2L where L is the hyperplane class, which generates the Picard group. The variety A parametrizing lines Y ⊂ X is an abelian surface [GH78, p. 779]. Four lines pass through a generic point x ∈ X [GH78, p. 781], so these lines dominate X. We have a(X, L) = a(Y, L|_{Y}) = 2 and b(X, L) = b(Y, L|_{Y}) = 1 so X is not balanced.

Suppose that X is defined over a number field F with X(F) Zariski dense; fix a metrization L of L. Manin’s formalism predicts the existence of an open set X^{∘} ⊂ X such that

#{x ∈ X^{∘}(F) : H_{L}(x) ≤ B} ∼ cB^{2}.

However, each line Y ⊂ X defined over F contributes

#{x ∈ Y (F) : H_{L}(x) ≤ B} ∼ c′(Y, L)B^{2},

where c′(Y, L) is a Tamagawa type number associated to Y and L. Moreover, after replacing F by a suitable finite extension these lines are Zariski dense in X, because rational points on abelian surfaces are potentially dense (see [HT00b, §3], for instance).

The weakly balanced property may fail too: Example 3.12. [BT96b] Let f, g be general cubic forms on ℙ^{3} and

X := {sf + tg = 0} ⊂ ℙ^{1} × ℙ^{3},

the Fano threefold obtained by blowing up the base locus of the pencil. The projection onto the first factor exhibits a cubic surface fibration

π : X → ℙ^{1},

Balanced line bundles 17

so that −K_{X} restricts to −K_{Y}, for every smooth fiber Y of π. Thus

a(Y, −K_{Y}) = a(X, −K_{X}) = 1.

Furthermore, the Néron-Severi rank of a smooth fiber of π is 7. On the other hand, by the Lefschetz theorem, we have rk NS(X) = 2 and

7 = b(Y, −K_{Y}) > b(X, −K_{X}) = 2,

i.e., −K_{X} is not weakly balanced on X.

Let Z be the union of singular fibers of π, exceptional curves in general smooth fibers Y, and the exceptional locus of the blow up to ℙ^{3}. Note that −K_{X} is balanced with respect to every rational curve on X which is not contained in Z.

4 del Pezzo surfaces

Let X be a smooth projective surface with ample −K_{X}, i.e., a del Pezzo surface. Their deformation types are determined by the degree of the canonical class d := (K_{X}, K_{X}). Basic examples are ℙ^{2} and ℙ^{1} × ℙ^{1}; more examples are obtained by blowing up 9 − d general points on ℙ^{2}. We have

• rk NS(X) = 10 − d;

• for 1 ≤ d ≤ 7 the cone Λ_{eff}(X) is generated by classes of exceptional curves, i.e., smooth rational curves of self-intersection −1.

Let L be a big Cartier divisor on X. When is it balanced? The only subvarieties of X on which we need to test the values of a and b are rational curves C ⊂ X, and b(C, L|_{C}) = 1.

It is easy to characterize curves breaking the balanced condition for the Fujita invariant. Let Z be the union of exceptional curves, if d > 1. Here, exceptional means in traditional sense. When d = 1, let Z be the union of exceptional curves plus singular rational curves in | − K_{X}|.

Lemma 4.1. Let X be a del Pezzo surface of degree d, C an irreducible rational curve with (C, C) ≠ −1, and L a big Cartier divisor on X. Then

a(C, L|_{C}) ≤ a(X, L),  (3)

i.e., L is weakly balanced on X outside of Z.

Proof. If C′ is an irreducible rational curve with (−K_{X}, C′) = 1 then C′ ⊆ Z. Indeed, if (C′, C′) < 0, then (C′, C′) = −1, by adjunction, and C′ is exceptional. On the other hand, if C′ and −K_{X} are linearly independent, the Hodge index theorem implies that d(C′, C′) − 1 < 0, i.e., (C′, C′) = −1 or 0. The second case is impossible since (K_{X}, C′) + (C′, C′) must be even. If C′ and −K_{X} are linearly dependent, then d(C′, C′) − 1 = 0 so that d = 1 and C′ is a singular rational curve in | − K_{X}|.

18 B. Hassett et al.

Let C′′ ⊂ X be a rational curve which is not in Z. After rescaling, we may assume that a(X, L) = 1, in particular, we do not assume that L is an integral divisor. Writing L + K_{X} ∼ D, where D is an effective ℚ-divisor, and computing the intersection with C′′ we obtain

(L, C′′) = (−K_{X}, C′′) + (D, C′′) ≥ (−K_{X}, C′′).

Since C′′ is not in Z, (−K_{X}, C′′) ≥ 2, i.e., (L, C′′) ≥ 2. It follows that

(L, C′′) + deg(KC̃′′) = (L, C′′) − 2 ≥ 0,

where C̃′′ is the normalization of C′′, i.e., a(C′′, L|_{C′′}) ≤ 1, as claimed.

We proceed with a characterization of b(X, L). Consider the Zariski decomposition

a(X, L)L + K_{X} = P + E,

where P is a nef ℚ-divisor and E = ∑_{i=1}^{n} e_{i}E_{i}, e_{i} ∈ ℚ_{>0}, (E_{i}, E_{j}) < 0. We have (P, E) = 0. Since X is a Mori dream space, P is semi-ample and defines a semi-ample fibration

π : X → B.

We have two cases:

Case 1. B is a point. Then

a(X, L)L + K_{X} = ∑_{i=1}^{n} e_{i}E_{i}

is rigid, and the classes E_{i} are linearly independent in NS(X). In particular, ⊕_{i}ℝ_{≥0}E_{i} is an extremal face of Λ_{eff}(X), and in fact the minimal extremal face containing a(X, L)L + K_{X}. It follows that

b(X, L) = rk NS(X) − n.

Case 2. B is a smooth rational curve. Then the minimal extremal face containing

a(X, L)L + K_{X} = P + ∑_{i=1}^{n} e_{i}E_{i}

is given by

NS_{π}(X) ∩ Λ_{eff}(X) = {P = 0} ∩ Λ_{eff}(X),

Balanced line bundles 19

where NS_{π}(X) ⊂ NS(X) is the subspace generated by vertical divisors, i.e., divisors D ⊂ X not dominating B. It follows that

b(X, L) = rk NS(X) − rk NS_{π}(X) = 1.

Proposition 4.2. Let X be a del Pezzo surface and L a big Cartier divisor on X. Then L is balanced if and only if a(X, L)L + K_{X} ∼ D, where D is a rigid effective divisor.

Proof. Assume that a(X, L) = 1. In Case 1, we must have

L + K_{X} = D = ∑_{i=1}^{n} e_{i}E_{i}, e_{i} > 0,

with E_{i} disjoint exceptional curves. Assume that L is not balanced so that b(X, L) = 1. Let π : X → ℙ^{2} be the blowdown of E_{1}, …, E_{n} and h a hyperplane class on ℙ^{2}. Then

L = −K_{X} + D = 3π^{∗}h + ∑_{i=1}^{n}(e_{i} − 1)E_{i}.

Let C be an irreducible rational curve which is not in Z. If C does not meet any of the E_{i} then

(L, C) = (3π^{∗}h, C) ≥ 3 > 2.

If C meets at least one of the E_{i} then

(L, C) = (−K_{X}, C) + (D, C) > 2

since the first summand is ≥ 2. It follows that a(C, L_{C}) < 1, i.e., L is balanced, contradicting our assumption.

In Case 2, we have

L + K_{X} = D = P + ∑_{i=1}^{n} e_{i}E_{i}, e_{i} ≥ 0,

where P is nef and E_{i} are disjoint exceptional divisors. Let π : X → ℙ^{1} be the fibration induced by the semi-ample line bundle P. The general fiber F of π is a conic and

rk NS(X) − rk NS_{π}(X) = 1.

We have (F, F) = 0, (−K_{X}, F) = 2, and the class of F is proportional to P. Hence, for any such F,

a(F, L|_{F}) = a(X, L), b(F, L|_{F}) = b(X, L) = 1.

Thus L is not balanced.

20 B. Hassett et al.

5 Equivariant geometry

Let G be a connected linear algebraic group, H ⊂ G a closed subgroup, and X a projective equivariant compactification of X^{∘} := H\G, a quasi-projective variety [Bor91, Ch. II]. Applying equivariant resolution of singularities we may assume that X is smooth and the boundary

∪_{α∈A}D_{α} = X \ X^{∘}

is a divisor with normal crossings with irreducible components D_{α}. If H is a parabolic subgroup of a semi-simple group G, then there is no boundary, i.e., A is empty, and H\G is a generalized flag variety which was discussed in Section 3. Throughout, we will assume that A is not empty.

Let 𝔛(G)^{∗} be the group of algebraic characters of G and

𝔛(G, H)^{∗} = { χ : G → 𝔾_{m} | χ(hg) = χ(g), ∀h ∈ H }

the subgroup of characters whose restrictions to H are trivial. Let Pic^{G}(X) be the group of isomorphism classes of G-linearized line bundles on X and Pic(X) the Picard group of X. For L ∈ Pic^{G}(X), the subgroup H ⊂ G acts linearly on the fiber L_{x} at x = H ∈ H\G. This defines a homomorphism

Pic^{G}(X) → 𝔛(H)^{∗}

to characters of H. Let Pic^{(G,H)}(X) be the kernel of this map. We will identify line bundles and divisors with their classes in Pic(X).

Proposition 5.1. Let G be a connected linear algebraic group and H a closed subgroup of G. Let X be a smooth projective equivariant compactification of X^{∘} := H\G with a boundary ∪_{α∈A}D_{α}. Then

1. we have an exact sequence

0 → 𝔛(G, H)^{∗} → ⊕_{α∈A}ℤD_{α} → Pic(X) → Pic(X^{∘}) → 0;

2. we have an exact sequence

0 → 𝔛(G, H)_{ℚ}^{∗} → Pic^{(G,H)}(X)_{ℚ} → Pic(X)_{ℚ};

and the last homomorphism is surjective when

ℭ(G, H) := Coker(𝔛(G)^{∗} → 𝔛(H)^{∗})

Balanced line bundles 21

or equivalently, Pic(X^{∘}), is finite.

3. we have a canonical injective homomorphism

Φ : ⊕_{α∈A}ℚD_{α} ↪ Pic^{(G,H)}(X)_{ℚ};

which is an isomorphism when ℭ(G, H) is finite.

Proof. The first statement is easy. The second assertion follows from [MFK94, Corollary 1.6] and [KKV89, Proposition 3.2(i)].

For the last assertion: Corollary 1.6 of [MFK94] implies that some multiple of D_{α} is G-linearizable. After replacing D_{α} by a multiple of D_{α}, we may assume that G acts on the finite-dimensional vector space H^{0}(X, O_{X}(D_{α})), via this G-linearization. Let s_{α} be the section corresponding to D_{α}. Then s_{α} ∈ H^{0}(X, O_{X}(D_{α})) is an eigenvector of the action by G. After multiplying by a character of G, if necessary, we may assume that s_{α} is fixed by the action of G. We let Φ(D_{α}) be this G-linearization.

Suppose that Φ(∑_{α} d_{α}D_{α}) = O_{X}, with trivial G-linearization, where d_{α} ∈ ℤ. Then there exists a rational function f such that

div(f) = ∑_{α} d_{α}D_{α}.

We may assume that f is a character of G whose restriction to H is trivial. By the definition of Φ, the function f must be fixed by the G-linearization. This implies that f ≡ 1. When ℭ(G, H) is finite, the surjectivity of Φ follows from (1) and (2).

From now on we consider the following situation: let H ⊂ M ⊂ G be connected linear algebraic groups. Typical examples arise when G is a unipotent group or a product of absolutely simple groups and H and M are arbitrary subgroups such that H\M is connected. Let X be a smooth projective G-equivariant compactification of H\G, and Y the induced compactification of H\M.

Lemma 5.2. Let π : X → X′ be a G-equivariant morphism onto a projective equivariant compactification of M\G. Assume that the projection G → M\G admits a rational section. Then

• π(D_{α}) = X′ if and only if D_{α} ∩ Y ≠ ∅;

• if D_{α} ∩ Y ≠ ∅ then D_{α} ∩ Y is irreducible;

• if D_{α} ∩ Y ≠ ∅ and D_{α′} ∩ Y ≠ ∅, for α ≠ α′ then D_{α} ∩ Y ≠ D_{α′} ∩ Y.

Proof. We have the diagram

22 B. Hassett et al.

\xymatrix{H\G \ar[d] & ⊂ & X \ar[d]_{π} & ⊃ & Y \ar[d]^{π} \\ M\G & ⊂ & X′ & ⊃ & M · e = point}

The first claim is evident. To prove the second assertion, choose a rational section σ : M\G −−→ G of the projection G → M\G. We may assume that a rational section is well-defined at a point M ∈ M\G. Consider the diagram

\xymatrix{D_{α} \ar[d]_{π} & ⊃ & D_{α}^{∘} \ar[d]^{π} \\ X′ & ⊃ & M\G}

where D_{α}^{∘} = D_{α} ∩ π^{−1}(M\G). We define a rational map

\begin{matrix}Ψ : D_{α}^{∘} & −−→ & D_{α} ∩ Y \\ x & ↦ & x · (σ ∘ π(x))^{−1}. \end{matrix}

Since Ψ is dominant, D_{α} ∩ Y is irreducible. Since the G-orbit of D_{α} ∩ Y is D_{α}^{∘}, the third claim follows.

Remark 5.3. When M is a connected solvable group, then G is birationally isomorphic to M × (M\G) so that the projection G → M\G has a rational section. See [Bor91, Corollary 15.8].

Theorem 5.4. Let

H ⊂ M ⊂ G

be connected linear algebraic groups. Let X be a smooth projective G-equivariant compactification of H\G and Y ⊂ X the induced compactification of H\M. Let L be a big Cartier divisor on X. Assume that

• the projection G → M\G admits a rational section;

• and a(X, L)L + K_{X} is linearly equivalent to a rigid effective ℚ-divisor D.

Furthermore, assume that either

1. Λ_{eff}(X) is finitely generated by effective divisors; or

2. there exists a birational contraction map f : X −−→ Z contracting D, where Z is a normal projective variety.

Then L is balanced with respect to Y.

Proof. Let X′ be any smooth projective equivariant compactification of M\G. We consider a G-rational map π : X −−→ X′ mapping

π : G ∋ g ↦ Mg ∈ M\G.

Balanced line bundles 23

After applying a G-equivariant resolution of the indeterminacy of the projection π if necessary, we may assume that π is a surjective morphism and X is a smooth equivariant compactification of H\G with a boundary divisor ∪_{α}D_{α}. Note that Y is a general fiber of π so that Y is smooth. Write the rigid effective ℚ-divisor D = a(X, L)L + K_{X} by

D = a(X, L)L + K_{X} = ∑_{i=1}^{n} e_{i}E_{i},

where E_{i}’s are irreducible components of a(X, L)L + K_{X} and e_{i} ∈ ℚ_{>0}. Our goal is to show that

(a(Y, L|_{Y}), b(Y, L|_{Y})) < (a(X, L), b(X, L)).

Since the E_{i}’s are rigid effective divisors, they are boundary components. This implies that

a(X, L)L|_{Y} + K_{Y} = (a(X, L)L + K_{X})|_{Y} = ∑_{i=1}^{n} e_{i}E_{i}|_{Y} ∈ Λ_{eff}(Y).

It follows that

a(Y, L|_{Y}) ≤ a(X, L).

Assume that a(Y, L|_{Y}) = a(X, L) =: a. Let F be the minimal supported face of Λ_{eff}(X) containing D = aL + K_{X} = ∑ e_{i}E_{i} and V_{F} a vector subspace generated by F. Either condition (1) or (2) guarantees that F is generated by E_{i}’s so that

b(X, L) = rk NS(X) − n.

(See Proposition 2.10 and Example 2.12.) Let F′ be the minimal supported face of Λ_{eff}(Y) containing

D|_{Y} = aL|_{Y} + K_{Y} = ∑ e_{i}E_{i}|_{Y} .

Let V′ be a vector subspace generated by all components of E_{i} ∩ Y. Since F′ contains all components of E_{i} ∩ Y, we have b(Y, L|_{Y}) ≤ codim(V′). Consider the restriction map:

Φ : NS(X)/V_{F} → NS(Y)/V′.

It follows from [KKV89, Proposition 3.2(i)], Lemma 5.2, and the exact sequence (1) in Proposition 5.1 that Φ is surjective. On the other hand, π^{∗}NS(X′) is contained in the kernel of Φ, so Φ has the nontrivial kernel. We conclude that

b(Y, L|_{Y}) ≤ codim(V′) < b(X, L).

24 B. Hassett et al.

Remark 5.5. Conditions (1) or (2) can be replaced by the condition: the numerical dimension ν(D) is zero (see [Leh13] for definitions). Note that (2) implies that the numerical dimension ν(D) is zero.

Corollary 5.6. Let H ⊂ M ⊂ G be connected linear algebraic groups and X a smooth projective equivariant compactification of H\G. Let Y ⊂ X be the induced compactification of H\M. Assume that the projection G → M\G admits a rational section. Then −K_{X} is balanced with respect to Y.

Note that −K_{X} is necessarily big by [FZ13, Thm. 1.2].

The existence of rational sections is important, and the second statement in Lemma 5.2 is not true in general:

Example 5.7. Consider the standard action of PGL_{3} on ℙ^{2}. Let ℙ^{5} be the space of conics and consider

X^{∘} = {T = (C, [p_{1}, p_{2}, p_{3}]) ∈ ℙ^{5} × Hilb^{[3]}(ℙ^{2}) | T satisfies (∗)},
(∗) : C is smooth, p_{i}’s are distinct, and p_{i} ∈ C.

Let X be the Zariski closure of X^{∘}, it is the Hilbert scheme of conics with zero dimensional subschemes of length 3, and is a smooth equivariant compactification of a homogeneous space S_{3}\PGL_{3}. Consider a ℙ^{2}- fibration f : X → Hilb^{[3]}(ℙ^{2}), the fiber over a general point Z ∈ Hilb^{[3]}(ℙ^{2}) is a ℙ^{2}, parametrizing conics passing through Z. On f^{−1}(Z), degenerate conics correspond to two lines passing through Z; these form three boundary components l_{i} on ℙ^{2}. However, general points on these components are on the same PGL_{3}-orbit, so there exists an irreducible boundary divisor D ⊂ X such that D ∩ f^{−1}(Z) = l_{1} ∪ l_{2} ∪ l_{3}. In other words, there is a non-trivial monodromy action on l_{i}’s. However, the monodromy action on the Picard group is trivial, and the balanced property still holds with respect these fibers.

6 Toric varieties

Manin’s conjecture for toric varieties was settled by Batyrev and Tschinkel via harmonic analysis on the associated adele groups in [BT98a] and [BT96a]. Implicitly, [BT96a] established a version of the balanced property. Here, we will use MMP to determine balanced line bundles. We expect that these techniques would also be applicable to some non-equivariant varieties. We refer to [FS04] for details concerning toric Mori theory, though most of properties we use hold formally for Mori dream spaces.

We start by recalling basic facts regarding toric Mori theory, which also hold for all Mori dream spaces (see [HK00, Section 1]):

Proposition 6.1 (D-Minimal Model Program). Let X be a ℚ-factorial projective toric variety and D a ℚdivisor. Then the minimal model program with respect to D runs, i.e.,

1. for any extremal ray R of NE_{1}(X), there exists the contraction morphism φ_{R};

Balanced line bundles 25

2. for any small contraction φ_{R} of a D-negative extremal ray R, the D-flip ψ : X −−→ X^{+} exists;

3. any sequence of D-flips terminates in finite steps;

4. and every nef line bundle is semi-ample.

Proof. See Theorem 4.5, Theorem 4.8, Theorem 4.9, and Proposition 4.6 in [FS04].

Proposition 6.2 (Zariski decomposition). Let X be a ℚ-factorial projective toric variety and D a ℚ-effective divisor. Applying D-MMP we obtain a birational contraction map f : X −−→ X′, with nef proper transform D′ of D. Consider a common resolution:

\xymatrix{ & X̃ \ar[dl]_{μ} \ar[d]_{ν} \\ X \ar@{-->}[r]_{f} & X′}

Then

1. μ^{∗}D = ν^{∗}D′ + E, where E is a ν-exceptional effective ℚ-divisor;

2. the support of E contains all divisors contracted by f;

3. if g : X̃ → Y is the semi-ample fibration associated to ν^{∗}D′, then for any ν-exceptional effective Cartier

divisor E′, the natural map

O_{Y} → g_{∗}O(E′),

is an isomorphism.

Proof. The assertions (1) and (2) follow from the Negativity lemma (see [FS04, Lemma 4.10]). Also see [FS04, Theorem 5.4].

The invariant b(X, L) can be characterized in terms of Zariski decomposition of a(X, L)L + K_{X}:

Proposition 6.3. Let X be a ℚ-factorial projective toric variety and D an effective ℚ-divisor on X. Suppose that

1. D = P + N, where P is a nef divisor and N ≥ 0;

2. let g : X → Y be the semi-ample fibration associated to P. For any effective Cartier divisor E which is

supported by Supp(N), the natural map

O_{Y} → g_{∗}O(E),

is an isomorphism.

26 B. Hassett et al.

Then the minimal extremal face of Λ_{eff}(X) containing D is generated by vertical divisors of g and components of N.

Proof. When D is big, the assertion is trivial. We may assume that dim(Y) < dim(X). Let F be the minimal extremal face of Λ_{eff}(X) containing D. Since F is extremal, it follows that F contains all vertical divisors of g and components of N.

On the other hand, our assumption implies that for general fiber X_{y}, N|_{X_{y}} is a rigid divisor on X_{y} (see [Leh14, Theorem 6.1]), and its irreducible components generate an extremal face F′ of Λ_{eff}(X_{y}). Let α ∈ \overline{NM}_{1}(X_{y}) be a nef cycle supporting F′ and F_{α} := \{α = 0\} ∩ Λ_{eff}(X). Since X_{y} is a general fiber, α ∈ \overline{NM}_{1}(X) so that F_{α} is an extremal face (see [Pet12, Theorem 6.8]). Since D · α = 0 and F is minimal we have F ⊂ F_{α}. Let D′ ∈ F be an effective ℚ-divisor. D′ is linear equivalent to a torus-invariant effective ℚ-divisor, so we may assume that D′ is a boundary divisor. Let T ⊂ Y be the big torus of Y. Then g^{−1}(T) ≅ X_{y} × T. Since D′ · α = 0, D′ is a sum of vertical divisors of g and components of N; our assertion follows.

Remark 6.4. Note that since the cone of pseudo-effective divisors is finitely generated, every extremal face is a supported face.

Proposition 6.5. Let X be a projective toric variety and Y an equivariant compactification of a subtorus of codimension one (possibly singular). Let L be a big Cartier divisor on X. Then L is weakly balanced with respect to Y.

Proof. Let M be the class of O_{X}(Y). Applying an equivariant embedded resolution of singularities, if necessary, we may assume that X and Y are smooth or at least ℚ-factorial terminal. Due to a group action of a torus, Y is not rigid, so that

a(X, L)L|_{Y} + K_{Y} = (a(X, L)L + K_{X})|_{Y} + M|_{Y} ∈ Λ_{eff}(Y).

Note that a(X, L)L + K_{X} is an effective ℚ-divisor on X. Thus we have

a(Y, L|_{Y}) ≤ a(X, L).

Suppose that a(Y, L|_{Y}) = a(X, L) =: a. Let D = aL + K_{X} + Y and consider the Zariski decomposition of D:

\xymatrix{ &  & X̃ \ar[dl]_{μ} \ar[d]_{ν} & ⊃ Ỹ \\ D ⊂ & X \ar@{-->}[r]_{f} & X′ & ⊃ D′,}

where D′ is the strict transform of D, which is nef, and Ỹ is the strict transform of Y. We may assume that both X̃ and Ỹ are smooth. Let F be the minimal extremal face of Λ_{eff}(X̃) containing

aμ^{∗}L + KX̃ + Ỹ .

Balanced line bundles 27

Since aμ^{∗}L + KX̃ ∈ F, it follows that codim(F) ≤ b(X, L). Since X has only terminal singularities, we have

aμ^{∗}L + KX̃ + Ỹ = aμ^{∗}L + μ^{∗}K_{X} + ∑_{i} d_{i}E_{i} + Ỹ ,

where d_{i}’s are positive integers and E_{i}’s are μ-exceptional divisors. It follows that F is the minimal extremal face containing μ^{∗}D and all μ-exceptional divisors. Let g : X̃ → B be the semi-ample fibration associated to ν^{∗}D′. Note that dim(B) < dim(X̃) since D is not big. Proposition 6.3 implies that F is generated by all vertical divisors of g and all ν-exceptional divisors. We denote the vector space, generated by F, by V_{F}.

Let F′ be the minimal extremal face of Λ_{eff}(Ỹ) containing

aμ^{∗}L|Ỹ + KỸ = (aμ^{∗}L + KX̃ + M̃)|Ỹ ,

where M̃ be the class of OX̃(Ỹ). Then F′ is also the minimal extremal face containing μ^{∗}D|Ỹ and all components of (E_{i} ∩ Ỹ)’s so that F′ is the minimal extremal face containing ν^{∗}D′|Ỹ and all components of (G_{j} ∩ Ỹ)’s, where G_{j}’s are all ν-exceptional divisors. In particular, F′ contains all vertical divisors of g|Ỹ : Ỹ → H = g(Ỹ). Since a multiple of ν^{∗}D′ admits a section vanishing along Ỹ, H is a Weil divisor of B, which is a subtoric variety. Let V′ ⊂ NS(Ỹ) be a vector space generated by vertical divisors of g|Ỹ and components of (G_{j} ∩ Ỹ)’s. Then b(Y, L) ≤ codim(V′). Consider the following restriction map:

Φ : NS(X̃)/V_{F} → NS(Ỹ)/V′.

We claim that Φ is surjective. Let N be an irreducible component of the boundary divisor of Ỹ which dominates H. There exists an irreducible component N′ of the boundary divisor of X̃ such that N′ contains N. Since N meets with a general fiber of g|Ỹ, N′ also meets with a general fiber of g. This implies that N′ dominates B. As in the proof of Lemma 5.2,

N′ ∩ Ỹ = mN + (vertical divisors of g|Ỹ).

Our claim follows from this. Hence

b(Y, L) ≤ codim(V′) ≤ codim(V_{F}) ≤ b(X, L).

Proposition 6.6. Let X be a ℚ-factorial terminal projective toric variety and L a big Cartier divisor on X. Suppose that the positive part of Zariski decomposition of D := a(X, L)L + K_{X} is nontrivial. Then L is not balanced.

28 B. Hassett et al.

Proof. After blowing up, if necessary, we may assume that D itself admits a Zariski decomposition

D = P + N,

where P is a nef ℚ-divisor and N ≥ 0 is the negative part. Let g : X → Y be the semi-ample fibration associated to P. We consider a general fiber X_{y} of g. Since a(X, L)L|_{X_{y}} + K_{X_{y}} = N|_{X_{y}} is a rigid effective divisor, we conclude that a(X, L) = a(X_{y}, L|_{X_{y}}). Let V ⊂ NS(X) be the vector space generated by vertical divisors of g and components of N and V′ ⊂ NS(X_{y}) the vector space generated by components of N|_{X_{y}}. The restriction map

Φ : NS(X)/V → NS(X_{y})/V′,

is surjective, by Lemma 5.2. On the other hand, let T be the big torus of Y. Then the preimage g^{−1}(T) of T is a product of T and a general fiber X_{y}. It follows that Φ is injective. Thus we have

b(X, L) = b(X_{y}, L|_{X_{y}}).

Hence L is not balanced on X.

An alternative proof of Theorem 5.4 for toric varieties is provided below:

Proposition 6.7. Let X be a ℚ-factorial terminal projective toric variety, L a big Cartier divisor on X, and Y an equivariant compactification of a subtorus of codimension one (possibly singular). Suppose that the positive part of the Zariski decomposition of a(X, L)L + K_{X} is trivial. Then L is balanced with respect to Y.

Proof. We follow the notations in the proof of Proposition 6.5. We only need to explain why b(Y, L|_{Y}) < b(X, L), when a(Y, L|_{Y}) = a(X, L). Since aμ^{∗}L + KX̃ is rigid, it follows that

codim(V_{F}) < b(X, L).

Thus our assertion follows.

Corollary 6.8. Let X be a ℚ-factorial terminal projective toric variety. A big Cartier divisor L is balanced with respect to all toric subvarieties if and only if a(X, L)L + K_{X} is rigid.

In general, it is possible that for a codimension one subtoric variety Y ⊂ X, the rank of NS(Y) exceeds the rank of NS(X). However, the balanced property still can be verified:

Example 6.9. Consider the standard action of 𝔾_{m}^{3} = {(t_{0}, t_{1}, t_{2})} on ℙ^{3} by

(t_{0}, t_{1}, t_{2}) · (x_{0} : x_{1} : x_{2} : x_{3}) ↦ (t_{0}x_{0} : t_{1}x_{1} : t_{2}x_{2} : x_{3}).

Balanced line bundles 29

Consider the subtorus

M = {(t_{0}, t_{1}, (t_{0}t_{1})^{−1})} ⊂ 𝔾_{m}^{3},

and let S be the equivariant compactification of M defined by

x_{0}x_{1}x_{2} = x_{3}^{3}.

This is a singular cubic surface with three isolated singularities of type A_{2}. We denote them by p_{1}, p_{2}, p_{3} ∈ ℙ^{3}. Since they are fixed under the action of 𝔾_{m}^{3} on ℙ^{3}, the blow-up B := Bl_{p_{1},p_{2},p_{3}}(ℙ^{3}) is an equivariant compactification of 𝔾_{m}^{3}. Moreover, the closure S̃ of M in B is the minimal desingularization of S and the class of S̃ in Pic(B) is ample. Put X := B × ℙ^{1} and Y := S̃ × ℙ^{1}. We have a diagram

\xymatrix{X \ar[d]_{π} & Y \ar@{^{(}->}[l] \ar[d] \\ B & S̃, \ar@{^{(}->}[l]}

Then Y is a nef divisor, and we have

rk NS(Y) = 8 > rk NS(X) = 5.

However, the anticanonical class −K_{X} is still balanced with respect to Y since

a(Y, −K_{X}|_{Y}) = a(X, −K_{X}) = 1
b(Y, −K_{X}|_{Y}) = 1 < b(X, −K_{X}) = 5.

This shows that, in general, we cannot expect to control the subgroup of NS(X) generated by vertical divisors. In the proof of Proposition 6.5, we were able to control the quotient by this subgroup.

Acknowledgments. The authors would like to thank Brian Lehmann for answering our questions and making useful comments. The authors also would like to thank the referees for useful suggestions which helped us improve the exposition. The second author would like to thank Fedor Bogomolov and Ilya Karzhemanov for useful discussions.

Funding. The first author was supported by National Science Foundation Grant 0901645, 0968349, and 1148609 the third author was supported by National Science Foundation Grants 0901777, 0968318, and 1160859.

30 B. Hassett et al.

References

[Abr09] D. Abramovich. Birational geometry for number theorists. In Arithmetic geometry, volume 8 of Clay Math. Proc., pages 335–373. Amer. Math. Soc., Providence, RI, 2009.

[BCHM10] C. Birkar, P. Cascini, Chr. D. Hacon, and James McKernan. Existence of minimal models for varieties of log general type. J. Amer. Math. Soc., 23(2):405–468, 2010.

[BDPP13] Sébastien Boucksom, Jean-Pierre Demailly, Mihai Paun, and Thomas Peternell. The pseudo-effective cone of a compact Kähler manifold and varieties of negative Kodaira dimension. J. Algebraic Geom., 22(2):201–248, 2013.

[Beh06] Roya Beheshti. Lines on projective hypersurfaces. J. Reine Angew. Math., 592:1–21, 2006.

[BFJ09] Sébastien Boucksom, Charles Favre, and Mattias Jonsson. Differentiability of volumes of divisors and a problem of Teissier. J. Algebraic Geom., 18(2):279–308, 2009.

[BM90] V. V. Batyrev and Yu. I. Manin. Sur le nombre des points rationnels de hauteur borné des variétés algébriques. Math. Ann., 286(1-3):27–43, 1990.

[Bor91] Armand Borel. Linear algebraic groups, volume 126 of Graduate Texts in Mathematics. Springer-Verlag, New York, second edition, 1991.

[Bou04] Sébastien Boucksom. Divisorial Zariski decompositions on compact complex manifolds. Ann. Sci. École Norm. Sup. (4), 37(1):45–76, 2004.

[BT95] Victor V. Batyrev and Yuri Tschinkel. Rational points of bounded height on compactifications of anisotropic tori. Internat. Math. Res. Notices, (12):591–635, 1995.

[BT96a] V. Batyrev and Yu. Tschinkel. Height zeta functions of toric varieties. J. Math. Sci., 82(1):3220–3239, 1996. Algebraic geometry, 5.

[BT96b] V. V. Batyrev and Y. Tschinkel. Rational points on some Fano cubic bundles. C. R. Acad. Sci. Paris Sér. I Math., 323(1):41–46, 1996.

[BT98a] V. V. Batyrev and Y. Tschinkel. Manin’s conjecture for toric varieties. J. Algebraic Geom., 7(1):15– 53, 1998.

[BT98b] V. V. Batyrev and Y. Tschinkel. Tamagawa numbers of polarized algebraic varieties. Astérisque, (251):299–340, 1998. Nombre et répartition de points de hauteur bornée (Paris, 1996).

[BT99] F. A. Bogomolov and Yu. Tschinkel. On the density of rational points on elliptic fibrations. J. Reine Angew. Math., 511:87–93, 1999.

Balanced line bundles 31

[Cam04] F. Campana. Orbifolds, special varieties and classification theory. Ann. Inst. Fourier (Grenoble), 54(3):499–630, 2004.

[CLT02] A. Chambert-Loir and Y. Tschinkel. On the distribution of points of bounded height on equivariant compactifications of vector groups. Invent. Math., 148(2):421–452, 2002.

[Cut86] Steven D. Cutkosky. Zariski decomposition of divisors on algebraic varieties. Duke Math. J., 53(1):149–156, 1986.

[DC12] Gabriele Di Cerbo. On Fujita’s log spectrum conjecture, 2012. arXiv:1210.5234.

[Deb03] Olivier Debarre. Lines on smooth hypersurfaces, 2003. www.math.ens.fr/∼debarre/Lines hypersurfaces.pdf.

[FMT89] J. Franke, Y. I. Manin, and Y. Tschinkel. Rational points of bounded height on Fano varieties.

Invent. Math., 95(2):421–435, 1989.

[FS04] O. Fujino and H. Sato. Introduction to the toric Mori theory. Michigan Math. J., 52(3):649–665, 2004.

[Fuj87] T. Fujita. On polarized manifolds whose adjoint bundles are not semipositive. In Algebraic geometry, Sendai, 1985, volume 10 of Adv. Stud. Pure Math., pages 167–178. North-Holland, Amsterdam, 1987.

[Fuj92] T. Fujita. On Kodaira energy and adjoint reduction of polarized manifolds. Manuscripta Math., 76(1):59–84, 1992.

[Fuj96] T. Fujita. On Kodaira energy of polarized log varieties. J. Math. Soc. Japan, 48(1):1–12, 1996.

[Fuj97] T. Fujita. On Kodaira energy and adjoint reduction of polarized threefolds. Manuscripta Math., 94(2):211–229, 1997.

[FZ13] Baohua Fu and De-Qi Zhang. A characterization of compact complex tori via automorphism groups. Math. Ann., 2013.

[GH78] Phillip Griffiths and Joseph Harris. Principles of algebraic geometry. Wiley-Interscience [John Wiley

& Sons], New York, 1978. Pure and Applied Mathematics.

[GTBT11] A. Gorodnik, R. Takloo-Bighash, and Y. Tschinkel. Multiple mixing for adele groups and rational

points, 2011. arXiv:1112.5743.

[HK00] Yi Hu and Sean Keel. Mori dream spaces and GIT. Michigan Math. J., 48:331–348, 2000. Dedicated to William Fulton on the occasion of his 60th birthday.

[HT99] B. Hassett and Y. Tschinkel. Geometry of equivariant compactifications of G_{a}^{n}. Internat. Math. Res. Notices, (22):1211–1230, 1999.

[HT00a] J. Harris and Y. Tschinkel. Rational points on quartics. Duke Math. J., 104(3):477–500, 2000.

32 B. Hassett et al.

[HT00b] Brendan Hassett and Yuri Tschinkel. Abelian fibrations and rational points on symmetric products. Internat. J. Math., 11(9):1163–1176, 2000.

[KKLV89] Friedrich Knop, Hanspeter Kraft, Domingo Luna, and Thierry Vust. Local properties of algebraic group actions. In Algebraische Transformationsgruppen und Invariantentheorie, volume 13 of DMV Sem., pages 63–75. Birkhäuser, Basel, 1989.

[KKV89] Friedrich Knop, Hanspeter Kraft, and Thierry Vust. The Picard group of a G-variety. In Algebraische Transformationsgruppen und Invariantentheorie, volume 13 of DMV Sem., pages 77–87. Birkhäuser, Basel, 1989.

[KMM87] Yujiro Kawamata, Katsumi Matsuda, and Kenji Matsuki. Introduction to the minimal model problem. In Algebraic geometry, Sendai, 1985, volume 10 of Adv. Stud. Pure Math., pages 283–360. North-Holland, Amsterdam, 1987.

[KMM92] János Kollár, Yoichi Miyaoka, and Shigefumi Mori. Rationally connected varieties. J. Algebraic Geom., 1(3):429–448, 1992.

[Laz04a] Robert Lazarsfeld. Positivity in algebraic geometry. I, volume 48 of Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of Modern Surveys in Mathematics. Springer-Verlag, Berlin, 2004. Classical setting: line bundles and linear series.

[Laz04b] Robert Lazarsfeld. Positivity in algebraic geometry. II, volume 49 of Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of Modern Surveys in Mathematics. Springer-Verlag, Berlin, 2004. Positivity for vector bundles, and multiplier ideals.

[Leh12] Brian Lehmann. A cone theorem for nef curves. J. Algebraic Geom., 21(3):473–493, 2012.

[Leh13] Brian Lehmann. Comparing numerical dimensions. Algebra Number Theory, 7(5):1065–1100, 2013.

[Leh14] Brian Lehmann. On Eckl’s pseudo-effective reduction map. Trans. Amer. Math. Soc., 366(3):1525– 1549, 2014.

[LT10] J. M. Landsberg and Orsola Tommasi. On the Debarre-de Jong and Beheshti-Starr conjectures on hypersurfaces with too many lines. Michigan Math. J., 59(3):573–588, 2010.

[MFK94] D. Mumford, J. Fogarty, and F. Kirwan. Geometric invariant theory, volume 34 of Ergebnisse der Mathematik und ihrer Grenzgebiete (2) [Results in Mathematics and Related Areas (2)]. Springer-Verlag, Berlin, third edition, 1994.

[MU83] Shigeru Mukai and Hiroshi Umemura. Minimal rational threefolds. In Algebraic geometry (Tokyo/Kyoto, 1982), volume 1016 of Lecture Notes in Math., pages 490–518. Springer, Berlin,

Balanced line bundles 33

[Nak04] Noboru Nakayama. Zariski-decomposition and abundance, volume 14 of MSJ Memoirs. Mathematical Society of Japan, Tokyo, 2004.

[Pet12] Thomas Peternell. Varieties with generically nef tangent bundles. J. Eur. Math. Soc. (JEMS), 14(2):571–603, 2012.

[Pey95] E. Peyre. Hauteurs et mesures de Tamagawa sur les variétés de Fano. Duke Math. J., 79(1):101–218, 1995.

[Som86] A. J. Sommese. On the adjunction theoretic structure of projective varieties. In Complex analysis and algebraic geometry (Göttingen, 1985), volume 1194 of Lecture Notes in Math., pages 175–213. Springer, Berlin, 1986.

[Tsc09] Y. Tschinkel. Algebraic varieties with many rational points. In Arithmetic geometry, volume 8 of Clay Math. Proc., pages 243–334. Amer. Math. Soc., Providence, RI, 2009.